I am working on a take home test that is due tomorrow. I seem to be struggling with lateral area and surface areas. Can you please help me understand.
The lateral area of a cone is about 226.2 square meters. If the slant height of the cone is 12 meters, what is the surface area to the nearest tenth.
Please help I have take home test due tomorrow and this is where i am struggling.
Tracy

Answer by mananth(16946)

The lateral area of a cone is about 226.2 square meters. If the slant height of the cone is 12 meters, what is the surface area to the nearest tenth.
Lateral area = pi*radius * slant height
226.2=pi*r*12
226.2/12*pi=r
r= 59.22 meters
..
Total area = lateral area + surface area of base
Area of base = pi*r^2
=pi*(59.22)^2
=11,017.6 m^2
Add up both the areas to get total surface area.
226.2+11,017
...
m.ananth@hotmail.ca